# SelectLLM: Query-Aware Efficient Selection Algorithm for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have gained increased popularity due to their remarkable success across various tasks, which has led to the active development of a large set of diverse LLMs. However, individual LLMs have limitations when applied to complex tasks because of such factors as training biases, model sizes, and the datasets used. A promising approach is to efficiently harness the diverse capabilities of LLMs to overcome these individual limitations. Towards this goal, we introduce a novel LLM selection algorithm called SelectLLM. This algorithm directs input queries to the most suitable subset of LLMs from a large pool, ensuring they collectively provide the correct response efficiently. SelectLLM uses a multi-label classifier, utilizing the classifier’s predictions and confidence scores to design optimal policies for selecting an optimal, query-aware, and lightweight subset of LLMs. Our findings show that the proposed model outperforms individual LLMs and achieves competitive performance compared to similarly sized, computationally expensive top-performing LLM subsets. Specifically, with a similarly sized top-performing LLM subset, we achieve a significant reduction in latency on two standard reasoning benchmarks: 13% lower latency for GSM8K and 70% lower latency for MMLU. Additionally, we conduct comprehensive analyses and ablation studies, which validate the robustness of the proposed model.
@@
Figure 1: Accuracy with majority vote accuracy for the test sets of GSM8K and MMLU against the number of LLMs considered from a diverse model pool. Models are added as per descending order of their performance on corresponding training sets.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in solving a wide range of core NLP tasks (Bommasani et al., 2021; Chang et al., 2023). Despite these advancements, it has been noted that existing LLMs struggle with complex tasks such as factually-grounded reasoning and planning (Wei et al., 2022; Kojima et al., 2022; Minaee et al., 2024). Moreover, the wide range of LLMs available seem to exhibit diverse capabilities (Jiang et al., 2023), resulting in no single (especially open-source) LLM being effective across all benchmarks and datasets.
@@
Though newer and larger models are being trained from scratch to close this gap, an alternative and cost-effective approach involves harnessing the diverse capabilities of existing models to improve the overall response quality using ensembling (Wang et al., 2022, 2023; Li et al., 2024) and collaborative frameworks (Wu et al., 2023b; Li et al., 2023). However, these approaches often require access to the responses from all models in the pool to choose the optimal response(s), which greatly increases the overall computational cost for such ensembles.
@@
Keeping in mind the diverse capabilities of LLMs, not all models may be apt for all kinds of tasks. Figure 1 reports the majority-vote accuracy of a model pool (spanning up to 7 diverse LLMs) on two challenging reasoning benchmarks – GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021). As the graph demonstrates, utilizing more LLMs initially improves performance, which is supported by previous research towards employing more LLMs and more responses per model (Li et al., 2024). However, note that using more (or even all) models in the pool does not necessarily result in the best scores. Thus, selectively abstaining from querying unsuitable LLMs for a given task may help improve the overall response quality of such ensembles. Additionally, such an approach would implicitly save computation by accessing fewer models per query.
In this paper, we propose the SelectLLM algorithm to explore this idea. Our approach first learns the model-specific capabilities for different LLMs through a dataset of diverse queries using a classifier. When performing inference for an unseen query, this insight is used to predict confidence scores for each LLM in the model pool, representing how likely they are to solve the underlying task. We create and test multiple selection policies that determine the optimal set of LLMs for the given query based on their respective confidence scores.
Our findings indicate that the proposed system: (i) Outperforms individual LLMs in terms of their accuracy and (ii) Achieves competitive accuracy score with a similar sized top-performing subset while significantly reducing latency costs. The contributions of our work are as follows:
@@
  • We introduce the novel SelectLLM algorithm, which is based on a multi-label classifier and an optimal confidence-based policy. This approach efficiently navigates input queries to the ideal subset of LLMs from a larger pool to improve response quality and also reduce computational costs in the process. To the best of our knowledge, this is the first study to propose LLM selection method.
  • The efficacy of the proposed algorithm is evaluated on two challenging reasoning benchmarks. We report a gain of $6.67$ points on GSM8K and $2.08$ points on MMLU in terms of accuracy compared to the highest scoring individual LLMs in the corresponding model pools. We also report significantly lower latencies (13% for GSM8K and 70% for MMLU) compared to similar sized top-performing LLM subset from the pools.
  • Through empirical analysis and ablation studies across two datasets and 7 LLMs, we demonstrate that the proposed model is reliable, robust, and cost-efficient.
@@
llama2-7b-lm 23.24 47.87
gemma-7b-lm 73.13 66.44
mistral-7b-lm 59.41 62.34
metamath-7b-lm 88.57 42.88
*gemma-7b-it 41.73 50.57
llama2-13b-chat 50.19 54.25
*mistral-7b-it 55.96 53.48
Table 1: Accuracy with majority voting (maj@10) for considered LLMs on GSM8K and MMLU datasets with all three splits. All scores were calculated over 10 response generations for each LLM. * The term ‘it’ indicates instruction-tuned LLMs.
@@ 2 Related Work
Model Diversity
Recent surveys (Bommasani et al., 2021; Minaee et al., 2024) suggest that LLMs can develop emergent capabilities. Specifically, this means that models can show behavior and demonstrate skills beyond those explicitly taught. By virtue of changing the training data, models can also be trained to exhibit a wide variety of domain expertise. At the same time, Jiang et al. (2023) demonstrate that no single open-source LLM outperforms other models across popular benchmarks. This further motivates the need to develop ensembling methods to improve the combined performance of a pool of LLMs with diverse capabilities.
Model Selection
Routing queries among LLMs involves aligning their capabilities with the underlying tasks from the input queries. However, selecting models for routing with LLMs differs from that in traditional ML (Bishop, 2006; Raschka, 2020) due to the disparity between the training and test datasets. LLMs are trained on massive corpora with straightforward objectives like next-token prediction (Brown et al., 2020), while test data often includes complex tasks like reasoning and question answering (Hendrycks et al., 2021; Cobbe et al., 2021; Joshi et al., 2017), summarization (Tam et al., 2023), and classification (Zhang et al., 2023), which may not be prevalent in the training data. This discrepancy makes it challenging to assess the difficulties in resolving complex queries. Additionally, studies like  Rabinovich et al. (2023) and  Srivatsa and Kochmar (2024) suggest that prompt characteristics such as length and readability can significantly impact an LLM’s ability to handle tasks.
@@
Previous attempts at ensembling and routing LLMs typically fall into two categories: (1) Selecting the best response from multiple LLM generations, as seen in Liu and Liu (2021), Ravaut et al. (2022), and Jiang et al. (2023). However, this approach requires querying all LLMs in the model pool for each query during inference, which can be computationally expensive with a large number of LLMs. (2) Minimizing the number of queries to larger LLMs to reduce latency and computational costs, as demonstrated by Shnitzer et al. (2023) and Ding et al. (2024), who redirect simpler queries to the smallest model capable of handling the task to minimize querying costs with minimal drop in performance. Tackling both tasks by routing to the single-best LLM has shown to be challenging (Srivatsa et al., 2024). In this paper, however, we aim to develop an algorithm that can improve response accuracy above individual LLMs and their combinations by querying only the subset of LLMs expected to be capable of solving the input query. This in turn, saves computational costs and reduces latency by avoiding unnecessary queries made to unsuitable LLMs.
@@ 3 Problem Setting
We propose an ensembling-based LLM inference algorithm – SelectLLM – to efficiently select the most suitable query-aware few LLMs from a large pool of available LLMs. The algorithm harnesses the diverse capabilities of different LLMs and selects a subset of models for the input query, jointly leading to the correct response and this selection of small subset leads to the reduction in latency.
Formally, for a given set of input queries $Q=\{q_{1},q_{2},\dots,q_{n}\}$ and a pool of LLMs $L=\{l_{1},l_{2},\dots,l_{k}\}$, the objective is to learn a selection model $M_{r}\in\{0,1\}^{k}$ (where 1 means that $L_{k}$ can answer the query correctly) such that each user query $q_{i}\in Q$ is directed to a subset $L_{s}$ with latency lower than the whole pool (i.e., $latency(L_{s})<latency(L)$). Here, $m_{i}$ is the $i^{th}$ predicted label by $M_{r}$, $L_{i}$ is the corresponding $i^{th}$ LLM in $L$, and $latency(L_{s})$ is the sum of the individual LLM latencies in $L_{s}$. $L_{s}$ consists of the top-$s$ LLMs based on the confidence scores predicted by $M_{r}$: in the best-case scenario, a query $q_{i}$ is processed by a single LLM, while in the worst-case scenario, $q_{i}$ is processed by all LLMs in $L$. We find that this algorithm favors optimal performance, surpassing individual LLM capabilities by a significant margin. Furthermore, it achieves competitive accuracy compared to top-performing similar size LLM subsets and maintaining lower latency.
@@ 4 Methodology
This section discusses the methodology behind LLM sampling and formally introduces the SelectLLM algorithm.
@@ 4.1.1 Selection of Benchmarks and LLMs
Recent studies suggest that LLMs struggle with reasoning and mathematical tasks (Patel et al., 2021; Wu et al., 2023a). This study focuses on two challenging reasoning tasks: specifically, we use the GSM8K dataset by Cobbe et al. (2021) for mathematical reasoning and the MMLU dataset by Hendrycks et al. (2021) for natural language understanding/reasoning tasks. GSM8K consists of 8,792 diverse grade-school level math word problems (MWPs) in English, while MMLU contains 15k multiple-choice questions in English, spanning 57 subjects across STEM, humanities, social sciences, etc. The dataset statistics for all three splits are summarized in Appendix Table 5.
We have selected a diverse set of LLMs based on explicit and implicit criteria. The explicit criteria encompass performance on benchmarks, training methodologies, model specialization, and modes of operation (chat and non-chat), among others. Some of these diverse attributes are presented in the Appendix Table 5. The implicit criteria include factors such as diverse inference latencies (refer to Table 6) and prompting types (i.e., zero-shot vs. few-shot), among others. Further, we consider relatively small open-source LLMs because: (i) Experiments with these LLMs are suitable for an academic lab setup; (ii) This aligns with the research trend towards developing LLMs suitable for small mobile devices Abdin et al. (2024); and (iii) We hypothesize that the proposed modeling approach, which is LLM-agnostic, will work for smaller LLMs and be scalable to larger LLMs, although we leave the proof of this to future work.
Based on the recent findings about the appropriate usage of prompts (Sahoo et al., 2024) and from our own experiments, we have tailored few-shot chain-of-thought (COT) prompts for non-chat LLMs and zero-shot COT prompts for chat LLMs. More details on LLM sampling and answer extraction are available in the Appendix Section A.
@@ 4.1.2 Data Preparation for the SelectLLM Model: SLData
In this study, we evaluate the performance of each LLM by generating 10 responses for each input query to ensure reliable and replicable behavior of the proposed model. We employ Majority Voting Li et al. (2024) to assess whether a query is correctly answered by the LLM or not. Specifically, Majority Voting (maj@k $\in\{0,1\}$) determines whether the most frequent answer from an LLM matches the gold answer or not. The accuracy with maj@10 across all input prompts is reported in Table 1. In the rest of this paper, we consider only those LLMs for which the answer extracted viability scores are above 90% (see more details in Appendix Section A) to ensure response reliability, resulting in 6 acceptable LLMs for the GSM8K dataset and 7 for the MMLU dataset. We prepare the training dataset for SelectLLM by associating each input query with the LLM(s) whose majority vote answer (across 10 samples) matches the gold answer, i.e., maj@10 $=1$. Formally, the target label for a query prompt $q\in Q$ is given by $label\left(q\right)=\left\{l\>|\>l\in L,maj@10\left(q,l\right)=1\right\}$, where $L$ is the set of candidate LLMs and $Q$ is the set of question prompts from GSM8K or MMLU. We denote this dataset as SLData.
[…]
3:for each query $q_{i}$ in $\mathcal{Q}$ do
7:     Initialize an empty set for answer set $A_{i}\leftarrow\phi$
8:     for each $l_{j}$ in $L_{s}$ do
9:         Generate 10 responses: $a_{i}^{10}=l_{j}(q_{i})$triangleright Generate 10 responses with LLM $l_{j}$
10:         Find answer frequency $a_{i}^{f}=\{a_{i}^{k}:\text{countof}(a_{i}^{k})/\sqrt{c_{i}^{(s)j}}\,|\,\text{for }a_{i}^{k}\text{ in unique}(a_{i}^{10})\}$
11:         $A_{i}\leftarrow A_{i}\cup a_{i}^{f}$
12:     end for
13:     Return most frequent answer in $A_{i}$
14:end for
@@ 4.2 The Proposed SelectLLM Algorithm
Prior research (Ding et al., 2024) indicates that easy queries are correctly solved by smaller or general-purpose LLMs, whereas complex queries necessitate the use of specialized or larger LLMs. Conversely, there are rare queries that are incorrectly responded to by large LLMs but correctly responded to by small LLMs Nezhurina et al. (2024). Due to a lack of widely established query-to-LLM mappings, brute-force approaches are typically employed, querying every available LLM to obtain correct answers. As language models continue to advance rapidly, resulting in a large number of LLMs, such approaches become computationally inefficient and sometimes infeasible.
To address this challenge, it is essential to explore modeling approaches to harness the capabilities of different LLMs (i.e., the ability to respond correctly to input queries) jointly and efficiently. One promising approach involves identifying and directing input queries to the most suitable subset of LLMs from the large pool. This ensures that the query is accurately addressed while maintaining lower latency compared to running inference on all LLMs in the pool. Towards this objective, we introduce the SelectLLM algorithm, designed to select a tailored subset of LLMs, taking into account the nature of the query and enabling them to collaboratively provide correct responses efficiently. The SelectLLM algorithm comprises two primary components: (i) A multi-label classifier (MLC) which is fine-tuned with SLData dataset, and (ii) A selection policy, which utilizes MLC’s prediction confidence scores (i.e., the likelihood of an LLM responding correctly to the query) to determine the suitable subset of LLMs.
@@
We fine-tune the pre-trained RoBERTa model (Liu et al., 2019) using SLData with a multi-label classification objective. The fine-tuned model predicts a set of LLMs (i.e., LLMs’ identities) that are capable of solving the input query. Specifically, the model learns to predict the most suitable LLMs along with corresponding confidence scores $C$. Since the multi-label classification (MLC) model is based on an encoder-only architecture, the cost of MLC inference is negligible compared to running inference with large auto-regressive language models (Sun et al., 2019). We have achieved weighted F1 scores of 0.71 and 0.68 for the GSM8K and MMLU test datasets, respectively for MLC.
@@
In this section, we discuss how the confidence scores $C$ are utilized to select a suitable subset of LLMs $L_{s}$ for each input query. $C$ is defined as $\{c_{1},c_{2},\dots,c_{i},\dots c_{n}\}$ where $c_{i}$ is the confidence scores for the $i^{th}$ input query. Each $c_{i}$ is further represented as $\{c_{i}^{l_{1}},c_{i}^{l_{2}},\dots c_{i}^{l_{j}}\dots c_{i}^{l_{k}}\}$ where $c_{i}^{l_{j}}$ is the confidence score of the $j^{th}$ LLM for the $i^{th}$ query.
The performance of SelectLLM is determined by the selection policy used. For example, a greedy policy that always selects the LLM (or set of LLMs) with the highest confidence can be suboptimal. This is because another LLM (or set of LLMs) in the pool might have higher accuracy but may not be chosen due to slightly lower confidence. Additionally, when there are two subsets with similar confidence and accuracy, it is more efficient to select the subset with lower cumulative latency. Considering these aspects, we propose the following three optimal policies:
@@
    WeightedMaxConf: This policy begins by selecting the top-$s$ LLMs based on their high confidence scores, i.e., for a given query $q_{i}$, we denote the selected LLMs as $L_{s}^{q_{i}}$. Subsequently, we modify the frequency of answer values extracted from the responses of each selected LLM, which involves dividing the frequency of each value by the square root of the confidence score associated with the respective LLM. Finally, we collect all response values and their modified frequencies across the selected LLMs (frequencies are added for the same value). The value with the highest frequency after the modification is selected as the final response. The formal steps are presented in Algorithm 1. Intuitively, dividing by the square root of the confidence score aims to mitigate biased predictions by the classifier. This adjustment ensures fairer opportunities for each selected LLM to contribute to the majority voting process.
Across all three policies, in case of a conflict where two LLMs have similar confidence, the light-weight LLM (i.e., with lower latency) is preferred.

### v3

@@ Abstract
Large language models (LLMs) have been widely adopted due to their remarkable performance across various applications, driving the accelerated development of a large number of diverse models. However, these individual LLMs show limitations in generalization and performance on complex tasks due to inherent training biases, model size constraints, and the quality or diversity of pre-training datasets. A promising direction is to efficiently harness the diverse capabilities of LLMs to overcome these individual limitations. To address these limitations, we introduce a novel LLM selection algorithm called SelectLLM, which efficiently directs input queries to the most suitable subset of LLMs from a large pool, ensuring that the selected models collectively provide accurate responses. SelectLLM employs a multi-label classifier and policy based on the classifier’s predictions and confidence scores in selecting an optimal, query-aware, and lightweight subset of LLMs. Our findings indicate that the proposed model outperforms existing ensemble-based baselines and achieves competitive performance with similarly sized top-performing LLMs while maintaining efficiency. Specifically, it achieves a huge reduction in inference latency on two challenging reasoning benchmarks: 13% on GSM8K and 70% on MMLU, compared to the top-performing baseline. Also, we establish a theoretical upper bound by an Oracle with LLMs and perform an in-depth linguistic analysis to understand the performance gap between the Oracle and SelectLLM.
@@
Figure 1: Accuracy (using majority voting) versus the number of LLMs plot for the GSM8K and MMLU test sets. Models are added in descending order of their performance on their corresponding training sets.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in solving a wide range of core NLP tasks (Bommasani et al., 2021; Chang et al., 2023). Despite these advances, existing LLMs still struggle with complex tasks such as factually-grounded reasoning and planning (Wei et al., 2022; Kojima et al., 2022; Minaee et al., 2024). Moreover, the wide range of LLMs available seem to exhibit diverse capabilities (Jiang et al., 2023), resulting in no single (especially open-source) LLM being effective across all benchmarks and datasets. Aligning with recent research trends DeepSeek-AI et al. (2025); Mirzadeh et al. (2024), we focus on natural language understanding tasks grounded in reasoning and do not consider natural language generation tasks.
@@
Although newer and more powerful models are constantly introduced, an alternative and cost-effective approach involves harnessing the diverse capabilities of existing models to improve the overall response quality using ensembling (Wang et al., 2022, 2023; Li et al., 2024) and collaborative frameworks (Wu et al., 2023; Li et al., 2023). However, these approaches often require access to the responses from all models in the pool to choose the optimal response(s), which greatly increases the overall computational cost for such ensembles.
@@
The individual LLMs exhibit diverse capabilities, i.e., not all models may be suited for all kinds of tasks. Fig. 1 reports the accuracy of a LLM model pool (spanning up to 7 diverse LLMs) on two challenging reasoning benchmarks – GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021). As the plot demonstrates, utilizing more LLMs initially improves performance, which is supported by previous research towards employing more LLMs and more responses per model (Li et al., 2024). However, note that using more (or even all) models in the pool does not necessarily result in the best scores overall. Thus, selectively abstaining from querying unsuitable LLMs for a given task may help improve the overall response quality of such ensembles. Additionally, such an approach would implicitly save computational resources by accessing fewer models per query.
In this paper, we propose the novel SelectLLM algorithm to explore this idea. Our approach first employs a multi-label classifier to learn the LLM-specific capabilities using a dataset of diverse queries. When running inference for a unseen query, this knowledge is utilized to predict confidence scores for each LLM in the model pool, reflecting their likelihood of successfully solving the task. Next, we develop various selection policies to determine the optimal subset of LLMs for each query based on these confidence scores and predictions. Additionally, we establish a theoretical Oracle model’s upper bound that can be achieved collectively by all LLMs and perform a qualitative and quantitative linguistic analysis of the inputs to understand the performance gap between the Oracle and SelectLLM. The contributions of our work are as follows:
@@
  • We introduce the novel SelectLLM algorithm, which is based on a multi-label classifier and an optimal confidence-based policy. This approach efficiently navigates input queries to the ideal subset of LLMs from a larger pool to improve response quality and simultaneously reduce computational costs.
  • The efficacy of the proposed SelectLLM algorithm is evaluated on two challenging reasoning benchmarks. We report an improvement of 1.90 points on GSM8K and 4.89 points in terms of accuracy on MMLU compared to the existing strong ensemble-based baselines and competitive performance with a model, which consists of similarly sized top-performing LLM subsets. Additionally, we observe significantly lower inference latencies, with reductions of 13% for GSM8K and 70% for MMLU, compared to the top-performing baseline.
  • We present a theoretical upper bound established by an Oracle model, representing the maximum performance achievable collectively by all LLMs in our pool. Furthermore, we conduct a linguistic feature-based analysis of the inputs to understand the gains achieved by SelectLLM and the performance gap relative to the Oracle model.
@@
Train Test Train Test
llama2-7b-lm 22.21 24.01 46.85 48.10
gemma-7b-lm 73.49 71.27 66.10 63.73
mistral-7b-lm 59.33 60.50 61.69 61.57
metamath-7b-lm 92.35 67.25 42.16 41.46
*gemma-7b-it 41.62 42.23 50.61 50.72
llama2-13b-chat 50.37 49.20 54.40 52.94
*mistral-7b-it 55.92 56.71 53.39 53.92
Table 1: Accuracy with majority voting (maj@M) for considered LLMs on GSM8K and MMLU datasets with train and test splits. All scores were calculated over M response generations for each LLM. Here we use M=10 inspired by Li et al. (2024).
@@ 2 Related Work
LLM Diversity and Capabilities:
LLMs exhibit emergent capabilities, enabling them to perform beyond their explicit training objectives, with diversity in training data fostering broad domain expertise (Bommasani et al., 2021; Minaee et al., 2024). This diversity extends to architectural variations, multilingual proficiency, and adaptability across domains, making them highly effective for a wide range of tasks. Additionally, LLMs demonstrate strong generalization and reasoning abilities, enabling them to tackle complex challenges such as question answering, summarization, and classification (Hendrycks et al., 2021; Cobbe et al., 2021; Joshi et al., 2017; Tam et al., 2023; Zhang et al., 2023). Despite their strengths, no single open-source LLM consistently outperforms others across benchmarks (Jiang et al., 2023), highlighting the necessity of ensemble methods to leverage diverse model strengths. This work focuses on leveraging the diverse capabilities of LLMs in SelectLLM to address individual LLM limitations.
@@
Previous attempts at ensembling and routing LLMs typically fall into three categories: (1) Selecting the best response from multiple LLM generations, as seen in Liu and Liu (2021), Ravaut et al. (2022), and Jiang et al. (2023). However, this approach requires querying all LLMs in the model pool for each query during inference, which can be computationally expensive with a large number of LLMs. (2) Minimizing the number of queries to larger LLMs to reduce latency and computational costs, as demonstrated by Shnitzer et al. (2023) and Ding et al. (2024), who redirect simpler queries to the smallest model capable of handling the task. This minimizes querying costs with minimal performance drop. Routing to the single-best LLM while balancing both accuracy and efficiency has proven to be challenging (Srivatsa et al., 2024). (3) Using multiple LLMs in a multi-agent collaboration setting Tran et al. (2025), where specialized LLMs work together to solve complex problems. However, it remains unclear how collaboration functions when non-specialized models are involved. The proposed model lays the foundation for addressing this gap. Moreover, we aim to develop an algorithm that improves response accuracy beyond individual LLMs and their combinations by querying only a subset of LLMs expected to be capable of solving the given query. This, in turn, reduces computational costs and latency by avoiding unnecessary queries to unsuitable LLMs.
@@ 3 Problem Setting
We propose an ensembling-based LLM inference algorithm – SelectLLM – to efficiently select the most suitable query-aware few LLMs from a large pool of available LLMs. The algorithm harnesses the diverse capabilities of different LLMs and selects a subset of models for the input query, jointly leading to the correct response, and this selection of a small subset leads to the reduction in latency.
Formally, for a given set of input queries $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{n}\}$ and a pool of LLMs =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{k}\}$, the objective is to learn a selection model $\mathcal{M}$ that selects a subset of LLMs $\mathcal{L}_{s}\subseteq\mathcal{L}$, which jointly produce the correct answer for a given input query $q_{i}\in\mathcal{Q}$, such that the cumulative latency satisfies $\text{latency}(\mathcal{L}_{s})<\text{latency}(\mathcal{L})$. In the best-case scenario, a query $q_{i}$ is processed by a single LLM, while in the worst-case scenario, $q_{i}$ is processed by all LLMs in $\mathcal{L}$.
@@ 4.1.1 Selection of Benchmarks and LLMs
As discussed in Section 1, LLMs often struggle with reasoning tasks. To advance modeling and performance in this domain, we have selected two challenging benchmarks. The GSM8K dataset, introduced by Cobbe et al. (2021), contains 8,792 grade-school level math word problems (MWPs) in English, focusing on mathematical reasoning. The second benchmark, MMLU, proposed by Hendrycks et al. (2021), includes 14,572 multiple-choice questions across 57 subjects, assessing multi-domain natural language understanding and reasoning. See Appendix Table 6 for detailed dataset statistics.
We have selected a diverse and sparse set of LLMs based on explicit and implicit criteria. The explicit criteria includes performance on benchmarks, training methodologies, model specialization, and modes of operation (chat vs. non-chat), among others. Some of these diverse attributes are presented in the Appendix Table 6. The implicit criteria include factors such as diverse inference latencies (refer to Table 7) and prompting types (i.e., zero-shot vs. few-shot), among others. Further, we consider relatively small open-source LLMs (yet representative of exiting open-source LLM space) because: (i) Experiments with these LLMs are suitable for an academic lab setup, and (ii) This aligns with the research trend towards developing LLMs suitable for small mobile devices Abdin et al. (2024). Furthermore, since these selection criteria ensure wide representativeness of the LLM pool, we hypothesize that if the proposed approach works for this pool of LLMs, it should be LLM pool-agnostic, although we leave the proof of this to future work. See Appendix A for more details.
@@ 4.1.2 Data Preparation for the SelectLLM Model: SLData
In this study, we evaluate the performance of each LLM by generating $M$ responses from it for each input query to ensure reliable and replicable behavior of the proposed model. We employ Majority Voting Li et al. (2024) to assess whether a query is correctly answered by the LLM or not. Specifically, Majority Voting (maj@M $\in\{0,1\}$) determines whether the most frequent answer from an LLM matches the gold answer or not. The accuracy with maj@M across all input prompts is reported in Table 1. In the rest of this paper, we consider only those LLMs for which the viable extracted answers are above 90% (see more details in Appendix Section A) to ensure response reliability, resulting in 6 acceptable LLMs for the GSM8K dataset and 7 for the MMLU dataset. We prepare the training dataset for the multi-label classification module of SelectLLM (detailed in the next section) by associating each input query with the LLM(s) whose majority vote answer (across $M$ samples) matches the gold answer, i.e., maj@M $=1$. Formally, the target label for a query prompt $q_{i}\in\mathcal{Q}$ is given by $label\left(q_{i}\right)=\left\{l\>|\>l\in\mathcal{L},maj@M\left(q_{i},l\right)=1\right\}$. We denote this dataset as SLData which is separately prepared for both GSM8K and MMLU.
@@ 4.2 The Proposed SelectLLM Algorithm
Building on the discussion in Section 1, prior research (Ding et al., 2024) indicates that easy queries are correctly solved by smaller or general-purpose LLMs, whereas complex queries necessitate the use of specialized or larger LLMs. Conversely, there are rare queries that are incorrectly responded to by large LLMs but correctly answered by smaller LLMs Nezhurina et al. (2024). Due to a lack of widely established query-to-LLM mappings, brute-force approaches are typically employed, querying every available LLM to obtain correct answers. As LLMs continue to advance rapidly in large numbers, such approaches become computationally inefficient and sometimes infeasible. One promising approach involves identifying and directing input queries to the most suitable subset of LLMs from a large pool, which jointly respond correctly. This ensures that the query is accurately addressed while maintaining lower latency compared to running inference on all LLMs in the pool.
Towards this objective, we introduce the query-aware SelectLLM algorithm, designed to select a tailored subset of LLMs, taking into account the nature of the query and enabling them to collaboratively provide correct responses efficiently. The SelectLLM algorithm comprises two primary components: (i) A multi-label classifier (MLC) which is fine-tuned with SLData dataset, and (ii) A selection policy, which utilizes MLC’s prediction and confidence scores (i.e., the likelihood of an LLM responding correctly to the query) to determine the suitable subset of LLMs.
@@
As indicated by prior studies Hu et al. (2024), lightweight language models such as BERT, RoBERTa, and T5 exhibit negligible lower query inference latency compared to LLMs. So, we developed a multi-label classifier (MLC) based on these models. The models were fine-tuned using SLData with a multi-label classification objective, incorporating label imbalance techniques Zhang et al. (2020) for both the GSM8K and MMLU datasets. The results are provided in Table 8. Among the tested models, the RoBERTa-based MLC outperformed the others, achieving weighted F1 scores of 0.71 and 0.68 for the GSM8K and MMLU test datasets, respectively. The fine-tuned model predicts a subset of LLMs (i.e., LLM identities) best suited to address the input query and confidence scores $C$ for each model in the pool.
@@
In this section, we discuss how the confidence scores $C$ are utilized to select a suitable subset of LLMs $\mathcal{L}_{s}$ for each input query. $C$ is defined as $\{c_{1},c_{2},\dots,c_{i},\dots c_{n}\}$ where $c_{i}$ is the confidence scores for the $i^{th}$ input query. Each $c_{i}$ is further represented as $\{c_{i}^{l_{1}},c_{i}^{l_{2}},\dots c_{i}^{l_{j}}\dots c_{i}^{l_{k}}\}$ where $c_{i}^{l_{j}}$ is the confidence score of the $j^{th}$ LLM for the $i^{th}$ query.
The performance of SelectLLM is determined by the selection policy used. For example, a greedy policy that always selects the LLM (or set of LLMs) with the highest confidence may be suboptimal. This is because another LLM (or set of LLMs) in the pool might have higher accuracy but may not be chosen due to slightly lower confidence. Additionally, when there are two subsets with similar confidence and accuracy, it is more efficient to select the subset with lower cumulative latency. Considering these aspects, we propose the following three optimal policies:
@@
    WeightedMaxConf: This policy begins by selecting the top-$s$ LLMs based on their high confidence scores, i.e., for a given query $q_{i}$, we denote the selected LLMs as $L_{s}^{q_{i}}$. Subsequently, we modify the frequency of answer values extracted from the responses of each selected LLM, which involves dividing the frequency of each value by the square root of the confidence score associated with the respective LLM. Finally, we collect all response values and their modified frequencies across the selected LLMs (frequencies are added for the same value). The value with the highest frequency after the modification is selected as the final response. The formal steps are presented in Algorithm 1 in Appendix E. Intuitively, dividing by the square root of the confidence score aims to mitigate biased selection of LLM by the policy similar to Wu et al. (2024). This adjustment ensures fairer opportunities for each selected LLM to contribute to the majority voting.
Across all three policies, in case of a conflict where two LLMs have similar confidence, the light-weight LLM (i.e., the lower latency) is preferred.